OpenStudy (anonymous):

On a certain sum, the compound interest accrued in the first two years is 5520 and that in the first three years is 9576. What is the rate of interest?

hartnn (hartnn):

@neta Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) Do you know this formula for compound interest ? \(\large F = P (1+r )^n \) you can use this, by putting n =3 first; then n=2, and just divide.

OpenStudy (anonymous):

f means

hartnn (hartnn):

oh sorry, i forgot to explain the terms , "F" means final amount after 'n' years, 'P' means principal amount, and 'r' means rate of interest in decimal.

hartnn (hartnn):

since you are given compound interest, say I, you can use, \(\large I = F-P = P (1+r )^n-P=P[(1+r^n)-1]\)

OpenStudy (anonymous):

k i will try

hartnn (hartnn):

good :) let me know if you couldn't do it.

OpenStudy (anonymous):

but here we dont know two things p and r

OpenStudy (anonymous):

i have consider two equations

hartnn (hartnn):

sorry for my typo, it should be \(\large I = F-P = P (1+r )^n-P=P[(1+r)^n-1]\) also, you'll have 2 equations, once when you put n=2 and then n=3 if you divide those ,P gets eliminated

OpenStudy (anonymous):

p(1+r)2-p(1+r)3=4056

hartnn (hartnn):

you have subtracted . try dividing the equations, P will get cancelled.

hartnn (hartnn):

and you'll just be left with an equation in r

OpenStudy (anonymous):

k

OpenStudy (anonymous):

but we have -p

hartnn (hartnn):

\(\huge \dfrac{I_1}{I_2}=\dfrac{P [(1+r)^{n_1}-1]}{P[(1+r)^{n_2}-1]}=\dfrac{ [(1+r)^{n_1}-1]}{[(1+r)^{n_2}-1]}\)

OpenStudy (anonymous):

r=1.734%

OpenStudy (anonymous):

i think i made mistake in divison

hartnn (hartnn):

yeah, try again...

OpenStudy (anonymous):

k thanks a lot

hartnn (hartnn):

welcome ^_^